Guile-GNOME: Clutter-GLX

version 1.10.0, updated 8 May 2012

Matthew Allum and OpenedHand LTD Intel Corporation

This manual is for (gnome clutter-glx) (version 1.10.0, updated 8 May 2012) Copyright 2006,2007,2008,2012 OpenedHand LTD Copyright 2009,2010,2011,2012 Intel Corporation

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. You may obtain a copy of the GNU Free Documentation License from the Free Software Foundation by visiting their Web site or by writing to:

The Free Software Foundation, Inc., 59 Temple Place - Suite 330, Boston, MA 02111-1307, USA

Short Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 X11 Specific Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 ClutterX11TexturePixmap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 GLX Specific Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5 Undocumented . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Chapter 1: Overview 1

1 Overview

(gnome clutter) wraps the Clutter graphical canvas toolkit for Guile. It is a part of Guile-GNOME.

See the documentation for (gnome gobject) for more information on Guile-GNOME.

2 X11 Specific Support

X11 specific API

2.1 Overview

The X11 backend for Clutter provides some specific API, allowing integration with the Xlibs API for embedding and manipulating the stage window, or for trapping X errors.

The ClutterX11 API is available since Clutter 0.6

2.2 Usage

[Function] clutter-x11-has-event-retrieval ⇒ (ret bool)

Queries the X11 backend to check if event collection has been disabled. ret TRUE if event retrival has been disabled. FALSE otherwise. Since 0.8

[Function] clutter-x11-get-stage-from-window (win unsigned-int32)
   ⇒ (ret <clutter-stage>)

Gets the stage for a particular X window. win an X Window ID ret A <clutter-stage>, or% NULL if a stage does not exist for the window. Since 0.8

[Function] clutter-x11-get-default-screen ⇒ (ret int)

Gets the number of the default X Screen object. ret the number of the default screen Since 0.6

[Function] clutter-x11-get-root-window ⇒ (ret unsigned-int32)

Retrieves the root window. ret the id of the root window Since 0.6

[Function] clutter-x11-get-stage-window (stage <clutter-stage>)
   ⇒ (ret unsigned-int32)

Gets the stages X Window.

stage  a <clutter-stage>

ret An XID for the stage window. Since 0.4

[Function] clutter-x11-set-stage-foreign (stage <clutter-stage>)
   (xwindow unsigned-int32) ⇒ (ret bool)

Target the <clutter-stage> to use an existing external X Window

stage  a <clutter-stage>

xwindow an existing X Window id

ret ‘#t’ if foreign window is valid

Since 0.4

[Function] clutter-x11-trap-x-errors

Traps every X error until clutter-x11-untrap-x-errors is called.

Since 0.6

[Function] clutter-x11-untrap-x-errors ⇒ (ret int)

Removes the X error trap and returns the current status.

ret the trapped error code, or 0 for success

Since 0.4

[Function] clutter-x11-has-composite-extension ⇒ (ret bool)

Retrieves whether Clutter is running on an X11 server with the XComposite extension

ret ‘#t’ if the XComposite extension is available

[Function] clutter-x11-set-use-argb-visual (use argb bool)

Sets whether the Clutter X11 backend should request ARGB visuals by default or not.

By default, Clutter requests RGB visuals.

If no ARGB visuals are found, the X11 backend will fall back to requesting a RGB visual instead.

ARGB visuals are required for the <"use-alpha"> property to work.

This function can only be called once, and before clutter-init is called.

use-argb ‘#t’ if ARGB visuals should be requested by default

Since 1.2

[Function] clutter-x11-get-use-argb-visual ⇒ (ret bool)

Retrieves whether the Clutter X11 backend is using ARGB visuals by default

ret ‘#t’ if ARGB visuals are queried by default

Since 1.2

[Function] clutter-x11-has-xinput ⇒ (ret bool)

Gets whether Clutter has XInput support.

ret ‘#t’ if Clutter was compiled with XInput support and XInput support is available at run time.

Since 0.8

[Function] clutter-x11-enable-xinput

Enables the use of the XInput extension if present on connected XServer and support built into Clutter. XInput allows for multiple pointing devices to be used. This function must be called before clutter-init. Since XInput might not be supported by the X server, you might want to use clutterx11-has-xinput to see if support was enabled. Since 0.8

[Function] clutter-x11-event-get-key-group (event <clutter-event>)

⇒ (ret int) Retrieves the group for the modifiers set in event

event  a  <clutter-event>  of  type  ‘CLUTTER_KEY_PRESS’  or
          ‘CLUTTER_KEY_RELEASE’

ret the group id

Since 1.4

3 ClutterX11TexturePixmap

A texture which displays the content of an X Pixmap.

3.1 Overview

<clutter-x11-texture-pixmap> is a class for displaying the content of an X Pixmap as a ClutterActor. Used together with the X Composite extension, it allows to display the content of X Windows inside Clutter.

The class uses the GLX EXT texture from pixmap OpenGL extension (http://people.freedesktop.org/~davidr/GLX EXT texture from pixmap.txt) if available

3.2 Usage

[Function] clutter-x11-texture-pixmap-new ⇒ (ret <clutter-actor>)

Creates a new <clutter-x11-texture-pixmap> which can be used to display the contents of an X11 Pixmap inside a Clutter scene graph

     ret  A new <clutter-x11-texture-pixmap>

Since 0.8

4 GLX Specific Support

GLX specific API

4.1 Overview

The GLX backend for Clutter provides some specific API for GLX related calls. The ClutterGLX API is available since Clutter 0.4

4.2 Usage

[Function] clutter-glx-texture-pixmap-new ⇒ (ret <clutter-actor>)

‘clutter_glx_texture_pixmap_new’ has been deprecated since version 1.6 and should not be used in newly-written code. Use clutter-x11-texture-pixmap-new instead

Creates a new, empty <clutter-glx-texture-pixmap>
ret  A new <clutter-glx-texture-pixmap>

Since 0.8

5 Undocumented

The following symbols, if any, have not been properly documented.

5.1 (gnome gw clutter-glx)

[Variable] clutter-x11-texture-pixmap-new-with-pixmap
[Variable] clutter-x11-texture-pixmap-new-with-window
[Variable] clutter-x11-texture-pixmap-set-automatic
[Variable] clutter-x11-texture-pixmap-set-pixmap
[Variable] clutter-x11-texture-pixmap-set-window
[Variable] clutter-x11-texture-pixmap-sync-window
[Variable] clutter-x11-texture-pixmap-update-area

Type Index 8

Type Index

(Index is nonexistent)

Function Index 9

Function Index

clutter-glx-texture-pixmap-new. . . . . . . . . . . . . . 6
clutter-x11-enable-xinput . . . . . . . . . . . . . . . . . . . 4
clutter-x11-event-get-key-group. . . . . . . . . . . . . 4
clutter-x11-get-default-screen. . . . . . . . . . . . . . 2
clutter-x11-get-root-window . . . . . . . . . . . . . . . . . 2
clutter-x11-get-stage-from-window . . . . . . . . . . 2
clutter-x11-get-stage-window . . . . . . . . . . . . . . . . 2
clutter-x11-get-use-argb-visual. . . . . . . . . . . . . 3

                                                clutter-x11-has-composite-extension . . . . . . . . 3
                                                clutter-x11-has-event-retrieval. . . . . . . . . . . . . 2

clutter-x11-has-xinput . . . . . . . . . . . . . . . . . . . . . . . 4

                                                clutter-x11-set-stage-foreign . . . . . . . . . . . . . . . 3
                                                clutter-x11-set-use-argb-visual. . . . . . . . . . . . . 3
                                                clutter-x11-texture-pixmap-new. . . . . . . . . . . . . . 5
                                                clutter-x11-trap-x-errors . . . . . . . . . . . . . . . . . . . 3
                                                clutter-x11-untrap-x-errors . . . . . . . . . . . . . . . . . 3